OpenStudy (anonymous):

Factor completely 4x2 – 8x – 60

OpenStudy (agreene):

\[4(x-5)(x+3)\]

OpenStudy (anonymous):

\[4x^2-8x-60=4(x^2-2x-15)=4(x-5)(x+3)\]